OpenStudy (jdolte):

I just need someone to check my answer! I'll fan and medal Use graphs and tables to find the limit and identify any vertical asymptotes of limit of 1 divided by the quantity x minus 6 as x approaches 6 from the left.

OpenStudy (jdolte):

Choices: -∞ ; x = 6 -∞ ; x = -6 ∞ ; x = -6 1 ; no vertical asymptotes

OpenStudy (jdolte):

I can't choose between A and C

OpenStudy (jdolte):

@phi @PartyShotz2409 @paki

OpenStudy (phi):

Approaching 6 "from the left" means "left of 6" i.e. from the side of 5 \[ \lim_{x\rightarrow 6^-} \frac{1}{x-6} \]

OpenStudy (phi):

if x is (for example) we get 1/(5-6) = 1/-1 = -1 as long as x is smaller than 6 we will get a minus number, right ? example: 5.9999 - 6 = -0.0001 and 1/ (-0.0001) = - 10000 (it looks like we are *very* negative as we get closer to 6)

OpenStudy (jdolte):

Ohh okay

OpenStudy (phi):

"limit" means we get close to x=6 (but never are allowed to reach it) it should be obvious the asympote is x=6 (and not x= -6, because we are not anywhere near x= -6)

OpenStudy (jdolte):

That makes sense! So it is A?

OpenStudy (phi):

yes

OpenStudy (jdolte):

Thanks! Can you help with a couple more pleaseee?

OpenStudy (phi):

one more

OpenStudy (jdolte):

Okay fair enough

OpenStudy (jdolte):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11).